A. Find the indicated sets.

2.ℙ({1,2,3,4})

6. ℙ({1,2})x ℙ({3})

10.  {Xϵ ℙ({1,2,3}):│X│≤1}


B. Suppose that │A│=m and │B│=n. Find the following cardinalities.

14. │ ℙ(ℙ(A))│

18. │ ℙ(Ax ℙ(B))│


1
Expert's answer
2021-03-18T07:25:55-0400

(2) Number of elements = 4

So power of set= 2n = 2n = 16

P({1,2,3,4})=16


(6) P({1,2}) =2n=4 and P({3}) =2n=2

P({1,2})×P({3})=4×2=8


(10)

P({1,2,3})=23=8


(14) |P(P(A))|= 4

P(A)= 21=2

P(P(A)) = 22=4



(18) P(B) = 21= 2

A× P(B) = 1 × 2 = 2

P(A×P(B))= 22= 4